OpenStudy (anonymous):

What property is shown? –8 + (–15) = (–15) + (–8) A. commutative property B. associative property C. opposite of a sum property D. distributive property

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@pooja195

OpenStudy (michele_laino):

hint: when I swap the operands, I'm applying the commutative property

OpenStudy (anonymous):

A. commutative property

OpenStudy (anonymous):

A? im not good with math

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

oh i pressed enter

OpenStudy (anonymous):

lol i have 4 more plz

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Which is an example of the opposite of a sum property? A. –17 + 32 = 32 + (–17) B. –26 + (4 + 8) = (–26 + 4) + 8 C. –5(6 + 3) = –5(6) + –5(3) D. –(–17 + 5) = 17 + (–5)

OpenStudy (anonymous):

i think D

OpenStudy (michele_laino):

in general the opposite of a sum is given by the subsequent formula: \[ - \left( {a + b} \right) = - a - b\] where a qnd b are two real numbers

OpenStudy (michele_laino):

and*

OpenStudy (michele_laino):

so, you are right!

OpenStudy (anonymous):

What is the value of the expression? 12.4 + (–2 + 1.6) A. 11 B. 12 C. 15 D. 16

OpenStudy (michele_laino):

we have to solve the computation inside the parentheses, so we have: 12.4+(-2+1.6)=12.4+(-0.4)=12.4-0.4=...?

OpenStudy (anonymous):

i think B

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (michele_laino):

at the rigt side we can write this: -3+(-14)=-3-14=...? please complete

OpenStudy (michele_laino):

right*

OpenStudy (anonymous):

-3+(-14)=-3-14=im thinking

OpenStudy (michele_laino):

ok! and what is -3-14=...?

OpenStudy (anonymous):

wait i think A

OpenStudy (anonymous):

am i correct

OpenStudy (michele_laino):

more precisely: right side= -17 now we can rewrite your original expression, as below: -(q+14)=-17 from which I can write this: q+14=17 so, what is q?

OpenStudy (anonymous):

-17

OpenStudy (michele_laino):

we have to find a number which added to 14 gives 17. what is such number?

OpenStudy (anonymous):

ok i had it but u r confusing me

OpenStudy (anonymous):

3 i guess

OpenStudy (michele_laino):

q=3 because 3+14=17

OpenStudy (anonymous):

ohhhh ok

OpenStudy (michele_laino):

that's right, your option is C

OpenStudy (anonymous):

Use an addition property to solve for c. 14 + (–12) = –12 + c A. –26 B. –14 C. 14 D. 26

OpenStudy (anonymous):

last 1

OpenStudy (michele_laino):

if I apply the commutative property, I can write this: 14+(-12)=-12+14 so, what is c?

OpenStudy (anonymous):

14=C

OpenStudy (michele_laino):

correct!!

OpenStudy (anonymous):

yay i got a 100

OpenStudy (michele_laino):

:) :)

OpenStudy (michele_laino):

congratulations!!